OpenStudy (anonymous):

2x + y = 5 x + 3y = 0

sam (.sam.):

Elimination method? 2x + y = 5 x + 3y = 0 Multiply by 2 ----------------------- 2x+y=5 2x+6y=0 -------- 0 -5y=0 y=0 Sub in equation 1, 2x=5 x=5/2 Sub in equation 2, x + 3(0)=0 x=0 ----------------------- When x=0, From equation 1, 2(0)+y=5 y=5